OpenStudy (anonymous):

MEDAL FOR ALLLLL+ ANY ANSWERS!!!! Which of the following ordered pairs is not a solution to the inequality y /> -3x - 2? A:(1, 1) B:(0, 0) C:(-1, 0) D:(0, -1)

OpenStudy (lgbasallote):

potato haha just kiddin what do you mean by />?

OpenStudy (anonymous):

\[\ge\]

OpenStudy (lgbasallote):

oh..well then try substituting each one is -3 (1) -2 /> 1 ? is -3(0) -2 /> 0? is -3(-1) -2 /> 0? -3(0) -2 /> -1 ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

idk man im taking classes online and they dont teach you nothing

OpenStudy (anonymous):

I know same here I take classes online too and they teach me nothing.

OpenStudy (anonymous):

ik someone plzz give me the answer cause its a test lol

OpenStudy (anonymous):

ok ill try lol

OpenStudy (anonymous):

i give you guys all medals when i get an answer that i can trust

OpenStudy (anonymous):

The possible answers that are in ordered pairs have two coordinates (x , y) for example (1, 1) the first 1 is the x-coordinate and the second 1 is the y-coordinate. You have to substitute the x and y's of each possible ordered pairs into the equation so: \[y \ge -3x - 2? \] let's try the first set of ordered pairs 1, 1 so you substitute 1 for the x and 1 for the y \[1 \ge -3 (1) - 2\] -3 times 1 is still -3 and -3 - 2 is -5. so -5 is not greater or equal to one. so it's not the first ordered pair. Now you just substitute the other pairs in to see if they are greater or equal to 1. :)

OpenStudy (anonymous):

*last sentence i meant "to see if they are greater or equal to your y-coordinate of each ordered pair. Sorry.

OpenStudy (anonymous):

just give me the answer because when i saw this question it was like chinese to me lol

OpenStudy (anonymous):

ok thnx

OpenStudy (anonymous):

i dont even know how to put it coordinates but thank you anyway

OpenStudy (anonymous):

actually the answer is B and D. xD